OpenStudy (anonymous):

Observe the graph and find how many zeros exist for this function: y = (2x – 2-x)

OpenStudy (sidsiddhartha):

to find zero of a function f(x) equate f(x)=0 then find x here y=(2x-2-x)=x-2 x-2=0 x=2 it is a zero

OpenStudy (anonymous):

its a linear function, so it only has one zero